# Structured coalescents, coagulation equations and multi-type branching processes

### v2

@@ 4.2. Characterization of the accumulation points
[…]
@@
The first expectation goes to 00 with $K\to\infty$ according to Lemma 4.6. For the second expectation, we find by Proposition 3.2
@@
5. Proofs of Theorems  2.4 and 2.5.
In this section, we will provide proofs of our main results. We will start this section with the proof of Theorem  2.4, which will follow as a combination of the results obtained in the previous sections. The proof of Theorem 2.5 will require some additional work due to the degenerate initial condition.
@@ 5.1. The critical sampling case
We know from Proposition 4.1 that the sequence $\left\{\mu^{K}\right\}_{K\geq 1}$ is tight. Furthermore, Proposition 4.10 tells us that its accumulation points are weak solutions of Eq. (2.3). Moreover, under Assumption 2.3 we have
$\mu^{K}_{i}(0,\{{\boldsymbol{n}}\})=\frac{L^{K}_{i}(0)}{K}\delta_{e_{i},{\boldsymbol{n}}}\xrightarrow[K\to\infty]{}c\beta_{i}\delta_{e_{i},{\boldsymbol{n}}}.$ (5.1)
Hence, the result follows from the uniqueness of solutions to Eq. (2.3) with initial condition (2.4) (see Proposition A.1 in Appendix  A). ∎
@@ 5.2. The large sampling case
We start this section with the last ingredients for the proof of Theorem 2.5, namely the convergence (in an appropriate sense) of the initial conditions of the empirical measures $({\boldsymbol{\mu}}^{K})_{K\geq 1}$. Recall that the problem arises in the large sampling case because
$\langle\mu_{i}^{K}(0),1\rangle=\frac{L_{i}^{K}(0)}{K}\xrightarrow[K\to\infty]{}\infty.$
@@
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\langle\mu_{i}^{K}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\,\rangle}\rangle=\lambda_{i}\beta_{i},\textrm{ in probability.}$.
@@
  • for each $\varepsilon\in(0,T)$, ${\boldsymbol{\mu}}^{K_{n}^{{}^{\prime\prime}}}\Rightarrow\boldsymbol{u}$ in $[\varepsilon,T]$.
@@
[…]
If the claim is true, then by Theorem 2.7 the function $\boldsymbol{u}$ would be uniquely determined (it is then in fact a deterministic measure-valued function) and would not depend on the choice of the initial subsequence $\left\{{\boldsymbol{\mu}}^{K_{n}}\right\}_{n\geq 1}$. The result would then follow.
@@
[…]
@@
Figure 4. An illustration of the two scenarios. Left: mono-chromatic scenario (configurations at time $\varepsilon/K$: $(5,0,0)$, $(11,0,0)$, $(4,0,0)$ in colony $1$, $(0,5,0)$, $(0,7,0)$, $(0,4,0)$, $(0,7,0)$ in colony $2$, $(0,0,7)$, $(0,0,9)$, $(0,0,7)$ in colony $3$). Right: poly-chromatic scenario (configurations at time $\varepsilon/K$: $(11,0,0)$, $(0,6,0)$, $(0,2,7)$ in colony $1$, $(5,7,0)$, $(0,0,3)$ in colony $2$, $(4,8,0)$, $(0,0,13)$ in colony $3$).
@@ 5.2.1. Mono-chromatic part
    Migrations of blocks to colony $i$ without any element of color $i$ do not affect $E_{i}^{K}$. Thus, when constructing $\hat{E}_{i}^{K}$ we can remove at time 00 all singletons with a color different from $i$ and only keep the singletons of color $i$.
@@ 5.2.2. Poly-chromatic part
The next result tells us that in colony $i$ the number of elements of a different color is much smaller than in $\gamma_{K}$.
@@ Appendix A Uniqueness and stochastic representation for the critical sampling
Before diving into the proof of Theorem  2.8, let us prove the following result about the uniqueness of solutions to the discrete coagulation equation.
@@
Assume that ${\boldsymbol{u}}$ is a solution of (2.3) with initial condition (2.4). Define ${\boldsymbol{\rho}}\coloneqq\langle{\boldsymbol{u}},1\rangle$. A straightforward calculation shows that ${\boldsymbol{\rho}}$ solves the system of equations
$\partial_{t}\rho_{i}=-\frac{\alpha_{i}}{2}\rho_{i}^{2}+\sum_{j\in[d]\setminus\{i\}}w_{j,i}\rho_{j}-w_{i,j}\rho_{i},\quad i\in[d],$ (A.1)
with initial condition ${\boldsymbol{\rho}}(0)={c}{\boldsymbol{\beta}}$. According to Picard-Lindelöf theorem, that initial value problem has at most solution. Since we have one solution, that solution is uniquely determined.
@@
$\partial_{t}u_{i}(t,{\boldsymbol{n}})=\alpha_{i}\left(\frac{1}{2}\sum_{{\boldsymbol{n}}_{1}+{\boldsymbol{n}}_{2}={\boldsymbol{n}}}u_{i}(t,{\boldsymbol{n}}_{1})u_{i}(t,{\boldsymbol{n}}_{2})-\rho_{i}(t)u_{i}(t,{\boldsymbol{n}})\right)+\ \sum_{j}w_{ji}$ $u_{j}(t,{\boldsymbol{n}})-w_{ij}u_{i}(t,{\boldsymbol{n}}),$
@@
$\sum_{{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}}\partial_{t}u_{i}(t,{\boldsymbol{n}}){\boldsymbol{\lambda}}^{\boldsymbol{n}}=$ $\alpha_{i}\bigg{(}\frac{1}{2}(\hat{u}_{i})^{2}-\langle u_{i},1\rangle\hat{u}_{i}\bigg{)}+\sum_{j\in[d]\setminus\{i\}}w_{j,i}\hat{u}_{j}-w_{i,j}\hat{u}_{i}$
@@
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=$ $c\alpha_{i}\beta_{i}\left(\frac{1}{2}v_{i}^{2}+\frac{1}{2}-v_{i}\right)\ +\ \sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}v_{j}w_{j,i}-v_{i}w_{i,j}-\sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}w_{j,i}-w_{i,j}.$ (A.4)
@@
Remark A.2 (Stochastic representation at equilibrium).
Instead of assuming that $d_{i}\geq 0$ for all $i\in[d]$, let us assume that ${\boldsymbol{\beta}}={\boldsymbol{\xi}}$ is the equilibrium probability measure, i.e.
We can then rewrite Eq. (A.4) as
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=$ $c\alpha_{i}\beta_{i}\left(\frac{1}{2}v_{i}^{2}+\frac{1}{2}-v_{i}\right)\ +\ \sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}v_{j}w_{j,i}-v_{i}w_{i,j}.$
Hence, proceeding as in the proof of Theorem 2.8, we deduce that
  • branches and dies at rate $\frac{{c}\alpha_{i}\beta_{i}}{2}$.
  • makes a transition from $i$ to $j$ at rate $\frac{\beta_{j}}{\beta_{i}}w_{j,i}$.

### v3

@@ 4.2. Characterization of the accumulation points
where $F_{i}({\boldsymbol{x}})=x_{i}$ is the projection to the $i$-th coordinate and ${\boldsymbol{f}}=(f,\dots,f)$, for $f\in C_{b}(\mathbb{R}_{+}^{d})$. Then
$\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{\infty})=0\text{, a.s.}$
@@
The first expectation goes to 0 with $K\to\infty$ according to Lemma 4.9. For the second expectation, we find by Proposition 3.2
@@
5. Proofs of Theorems 2.4 and 2.5.
In this section, we provide the proofs of our convergence results. We begin with the proof of Theorem 2.4, which follows directly from the results obtained in the previous sections. The proof of Theorem 2.5, on the other hand, requires additional work due to the degenerate initial condition.
@@ 5.1. The critical sampling case
Proposition 4.1 ensures that the sequence $\{\mu^{K}\}_{K\geq 1}$ is tight, while Proposition 4.10 identifies all of its accumulation points as weak solutions of Eq. (2.3). Together, these results already guarantee the existence of weak solutions to Eq. (2.3). In addition, under Assumption 2.3 we have
$\mu^{K}_{i}(0,\{{\boldsymbol{n}}\})=\frac{L^{K}_{i}(0)}{K}\delta_{e_{i},{\boldsymbol{n}}}\xrightarrow[K\to\infty]{}c\beta_{i}\delta_{e_{i},{\boldsymbol{n}}},$ (5.1)
which establishes convergence of the initial condition. The desired convergence thus follows from the uniqueness of solutions to Eq. (2.3) with initial condition (2.4) (see Prop. A.1 in Appendix A). ∎
@@ 5.2. The large sampling case
As mentioned at the beginning of this section, the proof of Theorem 2.5 requires an additional ingredient due to the degenerate initial condition, which comes from the fact that
$\langle\mu_{i}^{K}(0),1\rangle=\frac{L_{i}^{K}(0)}{K}\xrightarrow[K\to\infty$$\infty$.
@@
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\langle\mu_{i}^{K}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\,\rangle}\rangle=\lambda_{i}\beta_{i},\textrm{ in probability.}$
@@
  • for each $\varepsilon\in(0,T)$, ${\boldsymbol{\mu}}^{K_{n}^{{}^{\prime\prime}}}\Rightarrow\boldsymbol{u}$ in $\varepsilon,T]$.
@@
$\lim_{t\downarrow 0}\langle u_{i}(t),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle=\lambda_{i}\beta_{i}\text{, a.s.}$
If the claim holds, we can first conclude that $\boldsymbol{u}$ is a weak solution to (2.5) satisfying the initial condition (2.6). By Theorem 2.7, the function $\boldsymbol{u}$ is uniquely determined (and in fact becomes a deterministic measure-valued function), and therefore does not depend on the choice of the initial subsequence $\left\{{\boldsymbol{\mu}}^{K_{n}}\right\}_{n\geq 1}$. The result then follows.
@@
Taking $\limsup$ with $\varepsilon\downarrow 0$ on both sides, we find by Proposition 5.1 that
@@
Figure 5. An illustration of the two scenarios. Left: mono-chromatic scenario (configurations at time $\varepsilon/K$: $(5,0,0)$, $(11,0,0)$, $(4,0,0)$ in colony $1$, $(0,5,0)$, $(0,7,0)$, $(0,4,0)$, $(0,7,0)$ in colony $2$, $(0,0,7)$, $(0,0,9)$, $(0,0,7)$ in colony 33). Right: poly-chromatic scenario (configurations at time $\varepsilon/K$: $(11,0,0)$, $(0,6,0)$, $(0,2,7)$ in colony $1$, $(5,7,0)$, $(0,0,3)$ in colony $2$, $(4,8,0)$, $(0,0,13)$ in colony 33).
@@ 5.2.1. Mono-chromatic part
    Migrations of blocks to colony $i$ without any element of color $i$ do not affect $E_{i}^{K}$. Thus, when constructing $\hat{E}_{i}^{K}$ we can remove at time 0 all singletons with a color different from $i$ and only keep the singletons of color $i$.
@@ 5.2.2. Poly-chromatic part
The next result tells us that in colony $i$ the number of elements of a different color than $i$ is much smaller than $\gamma_{K}$.
@@ Appendix A Uniqueness and stochastic representation for the critical sampling
Before diving into the proof of Theorem 2.9, let us prove the following result about the uniqueness of solutions to the discrete coagulation equation.
@@
Assume that $\boldsymbol{u}$} is a solution of (2.3) satisfying the initial condition (2.4). Define ${\boldsymbol{\rho}}\coloneqq\langle{\boldsymbol{u}},1\rangle$. A straightforward calculation shows that $\boldsymbol{\rho}$} solves the system of equations
$\partial_{t}\rho_{i}=-\frac{\alpha_{i}}{2}\rho_{i}^{2}+\sum_{j\in[d]\setminus\{i\}}(w_{j,i}\rho_{j}-w_{i,j}\rho_{i}),\quad i\in[d],$ (A.1)
with initial condition ${\boldsymbol{\rho}}(0)={c}{\boldsymbol{\beta}}$. According to Picard-Lindelöf theorem, that initial value problem has at most one solution. Since we have one solution, that solution is uniquely determined.
@@
$\partial_{t}u_{i}(t,{\boldsymbol{n}})=\alpha_{i}\left(\frac{1}{2}\sum_{{\boldsymbol{n}}_{1}+{\boldsymbol{n}}_{2}={\boldsymbol{n}}}u_{i}(t,{\boldsymbol{n}}_{1})u_{i}(t,{\boldsymbol{n}}_{2})-\rho_{i}(t)u_{i}(t,{\boldsymbol{n}})\right)+\ \sum_{j}(w_{ji}$ $u_{j}(t,{\boldsymbol{n}})-w_{ij}u_{i}(t,{\boldsymbol{n}})),$
@@
$\sum_{{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}}\partial_{t}u_{i}(t,{\boldsymbol{n}}){\boldsymbol{\lambda}}^{\boldsymbol{n}}=$ $\alpha_{i}\bigg(\frac{1}{2}(\hat{u}_{i})^{2}-\langle u_{i},1\rangle\hat{u}_{i}\bigg)+\sum_{j\in[d]\setminus\{i\}}(w_{j,i}\hat{u}_{j}-w_{i,j}\hat{u}_{i})$
@@
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=$ $c\alpha_{i}\beta_{i}\left(\frac{1}{2}v_{i}^{2}+\frac{1}{2}-v_{i}\right)\ +\ \sum_{j\in[d]\setminus\{i\}}\Big(\frac{\beta_{j}}{\beta_{i}}v_{j}w_{j,i}-v_{i}w_{i,j}\Big)-\sum_{j\in[d]\setminus\{i\}}\Big(\frac{\beta_{j}}{\beta_{i}}w_{j,i}-w_{i,j}\Big).$ (A.4)